OpenStudy (anonymous):

Can someone please help me The functions of F and G are defined as follows f(x)=-3x^2-2 and g(x)=-4x-3 find f(3) and g(2) simplify the answer as much as possible

OpenStudy (anonymous):

so I know I do F(x) = -3x^2-2 F(3)=-3(3)^2-2(3) but im stuck after that what do i do now

OpenStudy (anonymous):

from there you simplify using order of operations. exponents → multiplication → subtraction

OpenStudy (anonymous):

where'd the second 3 come from in f(3)?

OpenStudy (anonymous):

so F(3)=-27-6 ?????

OpenStudy (anonymous):

sorry im really bad at math

OpenStudy (anonymous):

YOur original was f(x) = -3x² - 2, so f(3) = -3(3)² - 2. I think it should be f(3) = -27 - 2

OpenStudy (anonymous):

im so cofused lol how would i solve F(3)=-27-2 than ??

OpenStudy (anonymous):

-29 wouldnt be the answer would it ?

OpenStudy (anonymous):

yes. so it's -29. I was just saying that there shouldn't be an extra 3 at the end because there was no x there.

OpenStudy (anonymous):

okay thank you so than for g(x)=-4x-3 it would be g(2)=-4(2)-3 g(2)=-8-3 so than the answer would be -11 ?????

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

you're welcome